# DynamicPAE: Generating Scene-Aware Physical Adversarial Examples in Real-Time

### v1

@@ 4.1 Experiment Settings
Models Method AP50 AP01 SSIM LPIPS AP50 AP01 SSIM LPIPS
Yolov8-n Clean $0.864$ $0.921$ $0.933$ $0.941$
Noise $0.854_{\downarrow 1.0\%}$ $0.913_{\downarrow 0.8\%}$ 0.9892 0.0137 $0.905_{\downarrow 2.7\%}$ $0.918_{\downarrow 2.3\%}$ 0.9950 0.0070
PGD $\underline{0.577}_{\downarrow 28.7\%}$ $\underline{0.715}_{\downarrow 20.6\%}$ 0.9883 0.0218 $\underline{0.577}_{\downarrow 35.6\%}$ $\underline{0.644}_{\downarrow 29.7\%}$ 0.9947 0.0133
T-sea $0.604_{\downarrow 25.9\%}$ $0.750_{\downarrow 17.1\%}$ 0.9893 0.0158 $0.389_{\downarrow 54.3\%}$ $0.430_{\downarrow 51.1\%}$ 0.9953 0.0075
GAN-NAP $0.710_{\downarrow 15.4\%}$ $0.810_{\downarrow 11.0\%}$ $\underline{0.9897}$ $\underline{0.0127}$ $0.594_{\downarrow 33.9\%}$ $0.642_{\downarrow 30.0\%}$ 0.9954 0.0084
DynamicPAEλ=.8 $\mathbf{0.331}_{\downarrow 53.3\%}$ $\mathbf{0.438}_{\downarrow 48.3\%}$ 0.9912 0.0091 $\mathbf{0.307}_{\downarrow 62.6\%}$ $\mathbf{0.354}_{\downarrow 58.7\%}$ 0.9956 0.0069
Yolov5-s Clean $0.790$ $0.876$ $0.917$ $0.921$
Noise $0.757_{\downarrow 3.4\%}$ $0.848_{\downarrow 2.8\%}$ 0.9892 $\underline{0.0137}$ $0.851_{\downarrow 6.6\%}$ $0.865_{\downarrow 5.7\%}$ 0.9951 0.0068
PGD $\underline{0.270}_{\downarrow 52.1\%}$ $\underline{0.423}_{\downarrow 45.3\%}$ 0.9884 0.0223 $\underline{0.283}_{\downarrow 63.4\%}$ $\underline{0.375}_{\downarrow 54.6\%}$ 0.9948 0.0137
T-sea $0.482_{\downarrow 30.8\%}$ $0.631_{\downarrow 24.5\%}$ $\underline{0.9898}$ 0.0141 $0.365_{\downarrow 55.2\%}$ $0.410_{\downarrow 51.1\%}$ 0.9950 0.0091
GAN-NAP $0.401_{\downarrow 38.9\%}$ $0.564_{\downarrow 31.2\%}$ 0.9893 0.0148 $0.352_{\downarrow 56.5\%}$ $0.433_{\downarrow 48.8\%}$ $\underline{0.9952}$ 0.0086
DynamicPAEλ=.8 $\mathbf{0.225}_{\downarrow 56.6\%}$ $\mathbf{0.420}_{\downarrow 45.6\%}$ 0.9917 0.0075 $\mathbf{0.139}_{\downarrow 77.8\%}$ $\mathbf{0.256}_{\downarrow 66.5\%}$ 0.9955 0.0068
@@
Noise $0.840_{\downarrow 2.4\%}$ $0.892_{\downarrow 1.5\%}$ 0.9892 $\underline{0.0137}$ $0.837_{\downarrow 8.3\%}$ $0.848_{\downarrow 7.6\%}$ 0.9951 $\underline{0.0068}$
PGD $\underline{0.372}_{\downarrow 49.2\%}$ $\underline{0.488}_{\downarrow 41.8\%}$ 0.9884 0.0216 $0.364_{\downarrow 55.7\%}$ $0.420_{\downarrow 50.4\%}$ 0.9948 0.0130
T-sea $0.565_{\downarrow 29.9\%}$ $0.610_{\downarrow 29.7\%}$ $\underline{0.9900}$ 0.0139 $\underline{0.350}_{\downarrow 57.0\%}$ $\mathbf{0.372}_{\downarrow 55.1\%}$ 0.9953 0.0080
GAN-NAP $0.687_{\downarrow 17.7\%}$ $0.772_{\downarrow 13.5\%}$ 0.9896 0.0138 $0.517_{\downarrow 40.3\%}$ $0.558_{\downarrow 36.6\%}$ $\underline{0.9954}$ 0.0086
DynamicPAEλ=.8 $\mathbf{0.306}_{\downarrow 55.8\%}$ $\mathbf{0.427}_{\downarrow 47.9\%}$ 0.9918 0.0073 $\mathbf{0.323}_{\downarrow 59.7\%}$ $\underline{0.388}_{\downarrow 53.6\%}$ 0.9960 0.0049
@@
Noise $0.898_{\downarrow 2.1\%}$ $0.914_{\downarrow 1.8\%}$ 0.9891 0.0142 $0.895_{\downarrow 4.3\%}$ $0.898_{\downarrow 4.0\%}$ 0.9950 0.0072
PGD $\underline{0.323}_{\downarrow 59.6\%}$ $\underline{0.391}_{\downarrow 54.0\%}$ 0.9882 0.0219 $\underline{0.306}_{\downarrow 63.1\%}$ $\underline{0.338}_{\downarrow 60.1\%}$ 0.9946 0.0140
T-sea $0.679_{\downarrow 24.0\%}$ $0.728_{\downarrow 20.3\%}$ $\underline{0.9896}$ $\underline{0.0153}$ $0.774_{\downarrow 16.3\%}$ $0.784_{\downarrow 15.4\%}$ $\underline{0.9953}$ $\underline{0.0083}$
GAN-NAP $0.659_{\downarrow 26.0\%}$ $0.692_{\downarrow 23.9\%}$ 0.9893 0.0181 $0.640_{\downarrow 29.8\%}$ $0.663_{\downarrow 27.5\%}$ 0.9954 0.0108
DynamicPAEλ=.8 $\mathbf{0.116}_{\downarrow 80.3\%}$ $\mathbf{0.159}_{\downarrow 77.2\%}$ 0.9903 0.0118 $\mathbf{0.089}_{\downarrow 84.8\%}$ $\mathbf{0.124}_{\downarrow 81.4\%}$ 0.9950 0.0096
Faster- RCNN Clean $0.920$ $0.927$ $0.944$ $0.944$
Noise $0.909_{\downarrow 1.1\%}$ $0.916_{\downarrow 1.0\%}$ 0.9891 0.0138 $0.914_{\downarrow 3.0\%}$ $0.917_{\downarrow 2.7\%}$ 0.9951 $\underline{0.0068}$
PGD $\underline{0.478}_{\downarrow 44.2\%}$ $\underline{0.493}_{\downarrow 43.4\%}$ 0.9884 0.0201 $\underline{0.369}_{\downarrow 57.5\%}$ $\underline{0.378}_{\downarrow 56.6\%}$ 0.9948 0.0124
T-sea $0.831_{\downarrow 8.9\%}$ $0.844_{\downarrow 8.3\%}$ $\underline{0.9901}$ $\underline{0.0129}$ $0.601_{\downarrow 34.3\%}$ $0.614_{\downarrow 33.0\%}$ 0.9952 0.0097
GAN-NAP $0.753_{\downarrow 16.8\%}$ $0.768_{\downarrow 15.8\%}$ 0.9900 0.0144 $0.834_{\downarrow 11.0\%}$ $0.839_{\downarrow 10.5\%}$ 0.9958 0.0057
DynamicPAEλ=.8 $\mathbf{0.383}_{\downarrow 53.7\%}$ $\mathbf{0.398}_{\downarrow 52.9\%}$ 0.9911 0.0100 $\mathbf{0.235}_{\downarrow 70.9\%}$ $\mathbf{0.244}_{\downarrow 70.0\%}$ $\underline{0.9953}$ $0.0082$
The values bolded indicate the best results, the values underlined indicate the second-best results, the gray cells represent our method, and the subscript values represent the indicator decrease relative to the “Clean”.
@@ 4.2 Performances Evaluation
We evaluate performance based on settings similar to classical PAE studies [39] to focus on comparing PAE optimization or learning methods instead of the simulation. Furthermore, the stealthiness metrics are included to enhance the fairness for biased optimization.
We trained our model on the COCO dataset only and others on data identical to the evaluation datasets. The trade-off between the attack performance and the distortion of the original image is shown in Figure 5, demonstrating the significantly superior attack-distortion curve. Since our model can perform generation conditioned on ${\lambda}$ and automatically balance the attack and invisibility by the Skewness-Aligned Objective Re-weighting method, the curve in each sub-graph is evaluated in just one training session with the test-time settings of $\lambda=\{0.0,0.1,0.2,...,1.0\}$. However, the static PAE methods are only able to produce a single result, and thus our model achieves better practicality in PAE-driven model evaluation. The numerical results are shown in Table III, presenting the setting of $\lambda=0.8$ of DynamicPAE, which is generally more stealthy according to the attack-distortion curve. It is distinct that the dynamic PAE generated by our model achieves better performance, i.e., lower $AP_{50}$ and $AP_{01}$ values. Specifically, DynamicPAEλ=.8 boost the average drop of $AP_{50}$ to $-65.55\%$(1.95×1.95$\times$1.95 × on average) while preserving comparable stealthiness measured by SSIM, and only under one configuration (Yolov5-m, Inria) can T-sea achieve comparable aggressiveness. Furthermore, our model achieves 1.3×1.3$\times$1.3 × of $AP_{50}$ average drop compared to performing PGD on each image. In particular, for models with larger parameter sizes (Yolov3-m and Faster-RCNN), our approach exhibits more significant attack performance superiority, suggesting that our model can be scaled up for evaluating the adversarial robustness of larger-scale models. Furthermore, our method has not been trained on the Inria dataset, so the successful attacks on the Inria dataset indicate the zero-shot generalization capability in applying to proprietary data. We evaluate more transformation settings in the supplementary material and have the same result, except PGD performs better in zero spatial transformation.
We acknowledge that with a sufficiently large number of iterations, attacks based on traditional optimization, represented by PGD, under the same parameter distribution of $\boldsymbol{\oplus}$ and the same image $X$ may be able to achieve better attack results, but at an unacceptable time overhead. We present a comparison of time consumption in Table IV. All of the training and validation except DynamicPAEInf. is done on one Nvidia A40 with bf16 enabled. DynamicPAEVal., DynamicPAEInf. and PGD are timed through torch.cuda.Event after CUDA warmup, and they are tested on the Inria dataset with multiple runs. DynamicPAEVal. achieves a speedup of more than 2000 times compared to generating the PAE for each image with PGD and maintains higher attack performance. DynamicPAEInf. shows that our model can provide real-time patch generation even on personal graphics cards. Moreover, our training time is acceptable (about 40 hours on a single GPU). Evaluation of more PGD parameters is shown in section 4.3.1.
TABLE IV: Evaluation of GPU processing time. PGD and DynamicPAEVal. denote the average time of PAE generation on a batch of images with batchsize = 32, and DynamicPAEInf. shows the single-image inference latency. Static PAEs (T-sea and GAN-NAP) do not require inference computation, and thus they are not included.
@@
DynamicPAEInf. 11.910 ms (single image, 2080Ti)
4.3 On Overcoming the Noisy Gradient in PAE Training

### v3

@@ 4.1 Experiment Settings
Models Method AP50 AP01 SSIM LPIPS ASR↑ AP50 AP01 SSIM LPIPS ASR↑
Yolov8-n
Clean $0.864$ $0.921$ $0.933$ $0.941$
Noise $0.854_{\downarrow 1.0\%}$ $0.913_{\downarrow 0.8\%}$ 0.9892 0.0137 0.7% $0.905_{\downarrow 2.7\%}$ $0.918_{\downarrow 2.3\%}$ 0.9950 0.0070 2.6%
PGD $\underline{0.577}_{\downarrow 28.7\%}$ $\underline{0.715}_{\downarrow 20.6\%}$ 0.9883 0.0218 23.5% $\underline{0.577}_{\downarrow 35.6\%}$ $\underline{0.644}_{\downarrow 29.7\%}$ 0.9947 0.0133 25.0%
T-sea $0.604_{\downarrow 25.9\%}$ $0.750_{\downarrow 17.1\%}$ 0.9893 0.0158 22.7% $0.389_{\downarrow 54.3\%}$ $0.430_{\downarrow 51.1\%}$ 0.9953 0.0075 36.2%
GAN-NAP $0.710_{\downarrow 15.4\%}$ $0.810_{\downarrow 11.0\%}$ 0.9897 0.0127 11.9% $0.594_{\downarrow 33.9\%}$ $0.642_{\downarrow 30.0\%}$ 0.9954 0.0084 25.0%
CAP $0.581_{\downarrow 28.2\%}$ $0.724_{\downarrow 19.7\%}$ 0.9897 0.0148 25.6% $0.700_{\downarrow 23.3\%}$ $0.757_{\downarrow 18.5\%}$ 0.9956 0.0068 15.8%
DynamicPAEλ=.8 $\mathbf{0.331}_{\downarrow 53.3\%}$ $\mathbf{0.438}_{\downarrow 48.3\%}$ 0.9912 0.0091 42.0% $\mathbf{0.307}_{\downarrow 62.6\%}$ $\mathbf{0.354}_{\downarrow 58.7\%}$ 0.9956 0.0069 42.9%
Yolov5-s
Clean $0.790$ $0.876$ $0.917$ $0.921$
Noise $0.757_{\downarrow 3.4\%}$ $0.848_{\downarrow 2.8\%}$ 0.9892 0.0137 2.6% $0.851_{\downarrow 6.6\%}$ $0.865_{\downarrow 5.7\%}$ 0.9951 0.0068 3.9%
PGD $\underline{0.270}_{\downarrow 52.1\%}$ $\underline{0.423}_{\downarrow 45.3\%}$ 0.9884 0.0223 44.0% $\underline{0.283}_{\downarrow 63.4\%}$ $\underline{0.375}_{\downarrow 54.6\%}$ 0.9948 0.0137 41.3%
f T-sea $0.482_{\downarrow 30.8\%}$ $0.631_{\downarrow 24.5\%}$ 0.9898 0.0141 22.7% $0.365_{\downarrow 55.2\%}$ $0.410_{\downarrow 51.1\%}$ 0.9950 0.0091 30.3%
GAN-NAP $0.401_{\downarrow 38.9\%}$ $0.564_{\downarrow 31.2\%}$ 0.9893 0.0148 31.7% $0.352_{\downarrow 56.5\%}$ $0.433_{\downarrow 48.8\%}$ 0.9952 0.0086 29.5%
CAP $0.393_{\downarrow 39.8\%}$ $0.534_{\downarrow 34.2\%}$ 0.9899 0.0132 34.2% $0.653_{\downarrow 26.4\%}$ $0.695_{\downarrow 22.6\%}$ 0.9955 0.0072 13.6%
DynamicPAEλ=.8 $\mathbf{0.225}_{\downarrow 56.6\%}$ $\mathbf{0.420}_{\downarrow 45.6\%}$ 0.9917 0.0075 49.1% $\mathbf{0.139}_{\downarrow 77.8\%}$ $\mathbf{0.256}_{\downarrow 66.5\%}$ 0.9955 0.0068 55.4%
@@
Noise $0.840_{\downarrow 2.4\%}$ $0.892_{\downarrow 1.5\%}$ 0.9892 0.0137 2.1% $0.837_{\downarrow 8.3\%}$ $0.848_{\downarrow 7.6\%}$ 0.9951 0.0068 3.6%
PGD $\underline{0.372}_{\downarrow 49.2\%}$ $\underline{0.488}_{\downarrow 41.8\%}$ 0.9884 0.0216 37.6% $0.364_{\downarrow 55.7\%}$ $0.420_{\downarrow 50.4\%}$ 0.9948 0.0130 25.3%
T-sea $0.565_{\downarrow 29.9\%}$ $0.610_{\downarrow 29.7\%}$ 0.9900 0.0139 17.9% $\underline{0.350}_{\downarrow 57.0\%}$ $\mathbf{0.372}_{\downarrow 55.1\%}$ 0.9953 0.0080 26.6%
GAN-NAP $0.687_{\downarrow 17.7\%}$ $0.772_{\downarrow 13.5\%}$ 0.9896 0.0138 12.3% $0.517_{\downarrow 40.3\%}$ $0.558_{\downarrow 36.6\%}$ 0.9954 0.0086 19.2%
CAP $0.521_{\downarrow 34.4\%}$ $0.659_{\downarrow 24.7\%}$ 0.9899 0.0137 26.4% $0.545_{\downarrow 37.5\%}$ $0.583_{\downarrow 34.1\%}$ 0.9957 0.0060 13.8%
DynamicPAEλ=.8 $\mathbf{0.306}_{\downarrow 55.8\%}$ $\mathbf{0.427}_{\downarrow 47.9\%}$ 0.9918 0.0073 36.7% $\mathbf{0.323}_{\downarrow 59.7\%}$ $\underline{0.388}_{\downarrow 53.6\%}$ 0.9960 0.0049 28.1%
@@
Noise $0.898_{\downarrow 2.1\%}$ $0.914_{\downarrow 1.8\%}$ 0.9891 0.0142 1.5% $0.895_{\downarrow 4.3\%}$ $0.898_{\downarrow 4.0\%}$ 0.9950 0.0072 1.7%
PGD $\underline{0.323}_{\downarrow 59.6\%}$ $\underline{0.391}_{\downarrow 54.0\%}$ 0.9882 0.0219 39.6% $\underline{0.306}_{\downarrow 63.1\%}$ $\underline{0.338}_{\downarrow 60.1\%}$ 0.9946 0.0140 27.8%
T-sea $0.679_{\downarrow 24.0\%}$ $0.728_{\downarrow 20.3\%}$ 0.9896 0.0153 14.7% $0.774_{\downarrow 16.3\%}$ $0.784_{\downarrow 15.4\%}$ 0.9953 0.0083 5.2%
GAN-NAP $0.659_{\downarrow 26.0\%}$ $0.692_{\downarrow 23.9\%}$ 0.9893 0.0181 13.3% $0.640_{\downarrow 29.8\%}$ $0.663_{\downarrow 27.5\%}$ 0.9954 0.0108 16.0%
CAP $0.375_{\downarrow 54.4\%}$ $0.457_{\downarrow 47.4\%}$ 0.9889 0.0163 42.6% $0.464_{\downarrow 47.3\%}$ $0.500_{\downarrow 43.9\%}$ 0.9952 0.0085 23.1%
DynamicPAEλ=.8 $\mathbf{0.116}_{\downarrow 80.3\%}$ $\mathbf{0.159}_{\downarrow 77.2\%}$ 0.9903 0.0118 47.1% $\mathbf{0.089}_{\downarrow 84.8\%}$ $\mathbf{0.124}_{\downarrow 81.4\%}$ 0.9950 0.0096 45.3%
Faster-
RCNN
Clean $0.920$ $0.927$ $0.944$ $0.944$
Noise $0.909_{\downarrow 1.1\%}$ $0.916_{\downarrow 1.0\%}$ 0.9891 0.0138 0.5% $0.914_{\downarrow 3.0\%}$ $0.917_{\downarrow 2.7\%}$ 0.9951 0.0068 2.0%
PGD $\underline{0.478}_{\downarrow 44.2\%}$ $\underline{0.493}_{\downarrow 43.4\%}$ 0.9884 0.0201 13.0% $\underline{0.369}_{\downarrow 57.5\%}$ $\underline{0.378}_{\downarrow 56.6\%}$ 0.9948 0.0124 17.3%
T-sea $0.831_{\downarrow 8.9\%}$ $0.844_{\downarrow 8.3\%}$ 0.9901 0.0129 4.0% $0.601_{\downarrow 34.3\%}$ $0.614_{\downarrow 33.0\%}$ 0.9952 0.0097 15.5%
GAN-NAP $0.753_{\downarrow 16.8\%}$ $0.768_{\downarrow 15.8\%}$ 0.9900 0.0144 8.1% $0.834_{\downarrow 11.0\%}$ $0.839_{\downarrow 10.5\%}$ 0.9958 0.0057 7.9%
CAP $0.802_{\downarrow 11.9\%}$ $0.816_{\downarrow 11.1\%}$ 0.9900 0.0137 5.6% $0.861_{\downarrow 8.3\%}$ $0.865_{\downarrow 7.9\%}$ 0.9957 0.0061 3.6%
DynamicPAEλ=.8 $\mathbf{0.383}_{\downarrow 53.7\%}$ $\mathbf{0.398}_{\downarrow 52.9\%}$ 0.9911 0.0100 18.5% $\mathbf{0.235}_{\downarrow 70.9\%}$ $\mathbf{0.244}_{\downarrow 70.0\%}$ 0.9953 $0.0082$ 30.1%
RT-
DETR
-L
Clean $0.942$ $0.958$ $0.932$ $0.932$
Noise ${0.934}_{\downarrow 0.7\%}$ ${0.951}_{\downarrow 0.6\%}$ 0.9892 0.0137 0.6% ${0.905}_{\downarrow 2.7\%}$ ${0.912}_{\downarrow 2.0\%}$ 0.9950 0.0070 2.9%
PGD ${0.814}_{\downarrow 12.8\%}$ ${0.858}_{\downarrow 10.0\%}$ 0.9884 0.0223 9.4% ${0.750}_{\downarrow 18.2\%}$ ${0.763}_{\downarrow 16.9\%}$ 0.9948 0.0131 11.5%
T-sea ${0.933}_{\downarrow 0.9\%}$ ${0.951}_{\downarrow 0.7\%}$ 0.9888 0.0161 0.7% ${0.901}_{\downarrow 3.0\%}$ ${0.908}_{\downarrow 2.4\%}$ 0.9949 0.0092 3.5%
GAN-NAP ${0.933}_{\downarrow 0.9\%}$ ${0.950}_{\downarrow 0.8\%}$ 0.9897 0.0141 0.6% ${0.894}_{\downarrow 3.8\%}$ ${0.901}_{\downarrow 3.1\%}$ 0.9954 0.0090 3.6%
CAP $0.932_{\downarrow 1.0\%}$ $0.950_{\downarrow 0.8\%}$ 0.9894 0.0145 0.9% $0.908_{\downarrow 2.4\%}$ $0.914_{\downarrow 1.8\%}$ 0.9952 0.0076 2.8%
DynamicPAEλ=.8 $\underline{0.756}_{\downarrow 18.5\%}$ $\underline{0.815}_{\downarrow 14.3\%}$ 0.9915 0.0101 14.4% $\underline{0.617}_{\downarrow 31.5\%}$ $\underline{0.645}_{\downarrow 28.7\%}$ 0.9953 0.0090 19.8%
DynamicPAEλ=1. $\mathbf{0.573}_{\downarrow 36.9\%}$ $\mathbf{0.659}_{\downarrow 29.9\%}$ 0.9891 0.0164 31.2% $\mathbf{0.585}_{\downarrow 34.7\%}$ $\mathbf{0.615}_{\downarrow 31.7\%}$ 0.9951 0.0101 21.6%
The values bolded and underlined indicate the best and the second-best results, respectively. Subscripts show change relative to “Clean”.
Figure 8: Attack aggressiveness (measured by AP50, lower is better) & visual distortion (measured by LPIPS, higher is better) evaluated on COCO dataset, corresponding to Table V. Our framework DynamicPAE, with test-time settings $\lambda\in\{0.0,0.1,...,1.0\}$, significantly pushes the limit of aggressiveness-distortion trade-off.
@@ 4.2 Performances Evaluation
[…]
Note that the experiment is conducted on the white-box scenario to focus on comparing physical robustness and attack capability. The evaluation of black-box transfer attack, although DynamicPAE has not been designed for, is shown in the supplementary material. The results show that (1) when using the robust surrogate model, e.g., RT-DETR-L, the superiority of DynamicPAE is still maintained in the transfer attack scenario. (2) When using the non-robust light-weighted surrogate model, e.g., Yolo-v5s, T-sea may achieve slightly better transferability. The lower transferability of DynamicPAE in this case may be attributed to the higher model complexity of PAE. Given the greater superiority of (1) (ASR $18.4\%$) than (2) (ASR $6.2\%$), we believe DynamicPAE is also applicable to transfer attacks, and regularization may further improve the performance.
$\bullet$ Speed Evaluation: The total Floating Point Operations (FLOPs) and Parameters of the $\mathcal{G}$ that DynamicPAE framework successfully trained for attacking object detectors is only 5.45G and 29.29M, respectively, which ensures efficient inference on resource-limited edge devices. (The FLOPs of ResNet-50 is 4.09G). We present a comparison of time consumption in real devices in Table VI. All of the training and validation except DynamicPAEInf. is done on one Nvidia A40 with bf16 enabled. DynamicPAEVal., DynamicPAEInf. and PGD are timed through torch.cuda.Event after CUDA warmup, and they are tested on the Inria dataset with multiple runs. DynamicPAEVal. achieves a speedup of more than 2000 times compared to generating the PAE for each image with PGD and maintains higher attack performance. DynamicPAEInf. shows that our model can provide real-time patch generation even on personal graphics cards. Evaluation of more PGD parameters is shown in section 4.3.
TABLE VI: Evaluation of GPU processing time. PGD and DynamicPAEVal. denote the average time of PAE generation on a batch of images with batchsize = 32, and DynamicPAEInf. shows the single-image inference latency. Static PAEs (T-sea and GAN-NAP) do not require inference computation, and thus they are not included.
@@
DynamicPAEInf. 10.95 ms (Laptop RTX4050, 50W power limit)
TABLE VII: Comparison between training time and performance indicates the superiority of DynamicPAE in the overall adversarial evaluation efficiency.
T-sea GAN-NAP DynamicPAE
$AP_{50}$ Time $AP_{50}$ Time $AP_{50}$ Time
58.9% 0.7h 44.8% 0.2h 28.8% 0.7h
57.0% 1.3h 40.6% 0.5h 21.5% 1.4h
53.0% 2.7h 39.8% 1.0h 19.1% 2.8h
48.9% 5.4h 40.2% 2.0h 17.8% 5.7h
48.3% 10.8h 39.6% 3.9h 17.3% 11.3h
47.9% 21.6h 39.6% 7.8h 14.7% 22.6h
Considering the overall time consumption when evaluating the target model on the scenario defined by the task and the dataset, DynamicPAE still achieves superiority. We evaluate the training time when targeting Yolo-v5s on the COCO dataset, and show the performance of different frameworks with different training iterations compared to full training in Table VII. All training is done on a single A800 card, and the adaptive learning rate is performed by Adam. Although DynamicPAE needs to train the network, the time consumption is comparable, and its attack performance under the same training time is significantly higher. This is because both static attacks and the generative attacks in our framework require running backpropagation algorithms on the computational graph of the target model, whose overhead is comparable in magnitude to that of optimizing $\mathcal{G}$, as the FLOPs of evaluated detectors ranges from 8.7G (Yolo-v8n) to 267.37G (Faster-RCNN), which is much higher due to the high-resolution input settings of detectors. Also, the efficiency of our training framework benefits from the single forward-backward design shown in Algorithm 1.
4.3 On Overcoming the Noisy Feedback in Training